OpenStudy (anonymous):

How many times does the graph of y = 1 intersect the graph of y = x^2 + 1

OpenStudy (anonymous):

If you don't know this off the top of your head you could plug in 1 for y and solve for x. This is solving for the intersecting points.

OpenStudy (anonymous):

How far have you gotten in this problem?

OpenStudy (anonymous):

I dont understand it. Im retaking test from past modules and I realized i forgot alot of this stuff..

OpenStudy (anonymous):

y= 1 is a strait horizontal line the other function is a u shaped graph shifted up one I think it either tangent and it does touch once.

OpenStudy (anonymous):

two times try to draw grap then put x and y value y is strait line but x is parabola

OpenStudy (anonymous):

It hits once. at the minimum

OpenStudy (anonymous):

so it intersects twice?

OpenStudy (anonymous):

once

OpenStudy (anonymous):

tangent to the curve

OpenStudy (anonymous):

Solving by algebra: y=1 y=x^2+1 Set equal 1=x^2+1 0=x^2 x=0 plug in 0 for x y=1 (0,1) only point

OpenStudy (anonymous):